OpenStudy (anonymous):

Complete the order pairs so that each is a solution of the given linear equation. x=8-y where x is 0,7 and 8 find y

OpenStudy (campbell_st):

use substitution 0 = 8 - y can you find y..?

OpenStudy (anonymous):

is it -8?

OpenStudy (anonymous):

use y=8-x (on rearrangement and then put the values of x as 0,7,8 as 8,1,0) thus the pair is (0,8) (7,1) and (8,0)

OpenStudy (campbell_st):

so the 1st ordered pair is (0, -8) just repeat the process for x = 7 and x = 8

OpenStudy (anonymous):

(0,-8) (-7,1) (8,0)

OpenStudy (campbell_st):

and as @matricked has so kindly done... you don't to to gain an understanding... you can just copy his answers...

OpenStudy (anonymous):

correct?

OpenStudy (campbell_st):

not quite 7 = 8 - y whats y..?

OpenStudy (anonymous):

(7,1)

OpenStudy (campbell_st):

correct, lastly 8 = 8 - y find y..

OpenStudy (anonymous):

i mean (7,-1)

OpenStudy (campbell_st):

nope its just 1... 7 = 8 - 1 so y = 1

OpenStudy (anonymous):

@campbell_st i hope u mistyped the first ordered pair (0, -8)

OpenStudy (campbell_st):

thanks @matricked

OpenStudy (campbell_st):

and I did mistype the 1st solution its (0, 8) and the 2nd you found is (7, 1)